# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ 5.1.1 Quantile Estimation and Distribution Generation
We first assess the accuracy of the model’s quantile estimation. For each test observation $\mathbf{x}_{i}$, we predict $\tau_{j}$-quantiles, denoted as $\hat{y}_{ij}(\tau)$, and compare the empirical proportion of test responses that fall below these predicted quantiles, $\hat{\tau}_{j}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbb{I}(y_{i}\leq\hat{y}_{ij}(\tau))$, with the nominal level $\tau_{j}$, for $j=1,2,\cdots,m-1$. If the model perfectly estimates the $\tau_{j}$-quantile, $\hat{\tau}_{j}$ should be close to $\tau_{j}$, for $j=1,2,\cdots,m-1$. We conduct the experiment over 100 replications and plot the average $\hat{\tau}_{j}$ across these replications in Figure 5. The result demonstrates an excellent alignment between $\hat{\tau}_{j}$ and $\tau_{j}$ for all $j=1,2,\cdots,m-1$, confirming that the model provides accurate quantile estimates.
@@
Next, we evaluate the generative capabilities of the QRGMM model. To benchmark the generative performance, we conduct both unconditional and conditional tests, comparing the distributions of generated data by QRGMM and CWGAN model with the true test data.
@@
Using the trained QRGMM, we generate synthetic observations $\hat{y}^{\prime}_{i}$ for each test covariate vector $\mathbf{x}_{i}^{\prime}$, thus constructing a synthetic test dataset. Then we can compare the overall distribution of the generated samples $\{\hat{y}^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$ with the true test samples $\{y^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$. We conducted 100 experimental replications, calculating the sample means and standard deviations for the generated datasets and true test dataset in each run. The averages of these values across the 100 replications are reported in Table 1. To quantify distributional similarity, we also computed the Wasserstein Distance (WD) and the Kolmogorov-Smirnov (KS) statistic between the generated datasets and the true test dataset for each replication. The histograms of these metrics across replications are illustrated in Figure 6.
@@
The above comparison is unconditional, assessing overall distributional similarity. Since we know the true data-generating process in the synthetic setting, we can also perform a conditional test to evaluate the generative performance conditioned on a specific covariate vector. For a given covariate vector $\mathbf{x}$, we generate conditional datasets of $\hat{Y}(\mathbf{x})$ using QRGMM and CWGAN model and compare them with the true conditional dataset of $Y(\mathbf{x})$ sampled from the true model. Like the unconditional test, we perform 100 replications, calculating the sample means and standard deviations for each conditional dataset. The averaged results are also included in Table 1. The WD and KS values for the conditional test across replications are shown in Figure 7.
Table 1: Mean and Standard Deviation (SD) of Generated Samples.
@@
Truth $4349.4819$ $1433.0318$ $3948.5480$ $148.1465$
QRGMM $4348.6113$ $1433.1995$ $3944.1098$ $144.8040$
CWGAN $4325.2962$ $1376.1644$ $3939.9981$ $187.7608$
Figure 6: WD (left) and KS (right) for Unconditional Test.
Figure 7: WD (left) and KS (right) for Conditional Test.
[…]
@@
Figure 8: Histograms of generated vs. true data in one replication.
@@
Additionally, we visualize the generative performance by presenting histograms of the generated and true data distributions from one replication in Figure 8 for both the unconditional test and conditional test.
Summary of Generative Performance
@@
Our results demonstrate that QRGMM consistently achieves significantly lower WD and KS values compared to CWGAN in both unconditional and conditional tests (see Table 1 and Figures 67). This indicates that QRGMM provides a more accurate and robust reproduction of the true data-generating process, whereas CWGAN exhibits significant variability in performance across different replications. Furthermore, the histograms in Figure 8 illustrate that QRGMM-generated samples closely match the true data distribution in both unconditional and conditional settings, whereas CWGAN-generated samples exhibit noticeable discrepancies in conditional settings. These findings underscore the robustness and fidelity of QRGMM in modeling complex data distributions.
5.1.2 Risk Measures Approximation
@@
In this synthetic experiment, for a specified $\mathbf{x}$, the data-generating process of $Y(\mathbf{x})$ is known, enabling us to analytically compute true risk measures as benchmarks for assessing the performance of QRGMM and CWGAN approximations. In our setup, the loss function $L(l)$ defined in equation (1) is used, and the maximum loan amount $\bar{l}$ is set to the 99th percentile of $rY(\mathbf{x})$, with net unit revenue $r=1$. Loan levels $l$ are evaluated at 100 equally spaced points across $[2500,\bar{l}]$.
@@
We compute four true risk measures: $r_{1}(l)$ (probability of default), $r_{2}(l)$ (expected loss), and $\text{LGD}(l)$ (loss given default), as specified in Section 2, and $r_{3}(l)=\mathbb{E}[((l-rY(\mathbf{x}))^{+})^{2}]$, a specific form of the generalized risk measure highlighting severe losses. QRGMM and CWGAN approximations are calculated over 100 replications, each with 10,000 samples, and their means are compared to true values and estimations using the true test data samples averaged over 100 replications. These results are depicted in Figure 9, with subplots for $r_{1}(l)$, $r_{2}(l)$, $\text{LGD}(l)$, and $r_{3}(l)$.
Figure 9: Approximations of risk measures across loan levels $l$.
The results are highly encouraging: QRGMM’s approximations consistently track the true values across all measures, demonstrating robust accuracy. Conversely, CWGAN exhibits notable discrepancies, especially in $r_{2}(l)$ and $r_{3}(l)$, reflecting weaker performance. This comparison underscores QRGMM’s superior capability in capturing complex risk measures, QRGMM’s reliability significantly outshines CWGAN, demonstrating its applicability in SCF credit risk management. Notably, these figures vividly offer a clear, immediate view of how risk measures evolve with loan amount $l$, providing decision-makers with an intuitive tool for understanding the risk dynamics, thereby greatly facilitating loan sizing decisions.
@@ 5.2.1 Big Mart Sales Dataset
Figure 10: Alignment of Nominal and Empirical Quantile Levels on the Big Mart Data.
Unconditional Distribution Comparison.
Since the true data-generating mechanism is unknown in real-world datasets, we cannot perform a conditional test as in the synthetic experiment. Instead, we focus on an unconditional comparison of the generated and observed sales distributions. We trained the QRGMM and CWGAN models on the training set and subsequently generated a synthetic test dataset for each model across 100 replications.
Table 2 summarizes the means and standard deviations (averaged over the 100 replications) of both the generated and observed test sales. QRGMM shows a closer match to the true data than CWGAN, suggesting that QRGMM can capture the overall distribution of sales more accurately. To further quantify distributional similarity, we compute the WD and KS statistic for each replication, and plot their distributions in Figure 11. QRGMM again exhibits lower WD and KS values than CWGAN, reinforcing its superior generative ability in this real-world scenario.
Table 2: Mean and Standard Deviation of Generated Samples.
@@
      Truth       $2177.6046$       $1708.2316$
      QRGMM       $2168.4776$       $1680.5823$
      CWGAN       $2194.7852$       $1518.2655$
Figure 11: WD (left) and KS (right) for Big Mart Sales.
Figure 12 illustrates the observed and generated sales distributions in one representative replication. The QRGMM-generated distribution tracks the empirical distribution more closely than does CWGAN.
Figure 12: Histogram Comparison of Observed vs. Generated Big Mart Sales (Unconditional).
Discussion.
In summary, QRGMM continues to exhibit strong performance in quantile estimation and distribution generation on the Big Mart Sales dataset. Although a conditional test is infeasible due to unknown true mechanisms, the unconditional metrics and visual checks consistently show QRGMM outperforming CWGAN in capturing the real data distribution. This result underscores QRGMM’s practicality for complex SCF applications.
@@ 5.2.2 Amazon CBEC Dataset
Unconditional Distribution Comparison.
Since the true conditional distribution remains unknown, we again adopt an unconditional evaluation strategy. Both QRGMM and CWGAN are trained on the training set and used to generate synthetic test sales data over 100 replications.
Table 3 reports the means and standard deviations (averaged across 100 replications) of the generated and observed test sales. Similar to the results on Big Mart dataset, QRGMM aligns more closely with the true data than CWGAN. In addition, Figure 14 compares the WD and KS statistics across replications. QRGMM consistently yields lower values for both metrics, demonstrating its superior generative fidelity in this CBEC scenario.
Table 3: Mean and Standard Deviation of Generated Samples.
@@
      QRGMM       $39.2979$       $76.7818$
      CWGAN       $36.0173$       $60.7141$
Figure 14: WD (left) and KS (right) for Amazon CBEC Sales.
Figure 15 presents a histogram comparison of observed and generated sales distributions from one representative replication. As with the Big Mart dataset, the QRGMM-generated samples align more closely with the observed sales, whereas CWGAN shows noticeable deviation.
Figure 15: Histogram Comparison of Observed vs. Generated Amazon CBEC Sales (Unconditional).
Discussion.
Overall, QRGMM continues to demonstrate robust performance in both quantile estimation and unconditional distribution generation on the Amazon CBEC dataset. Despite the complex interaction effects among product and sales teams, as well as the influence of various marketing and product attributes, the model effectively captures the data structure. These results affirm QRGMM’s practical utility in complex CBEC environments, enhancing its relevance for SCF risk management.

### v3

@@ 6.1.1 Quantile Estimation and Distribution Generation
We first assess the accuracy of the model’s quantile estimation. For each test observation $\mathbf{x}_{i}$, we predict $\tau_{j}$-quantiles, denoted as $\hat{y}_{ij}$, and compare the empirical proportion of test responses that fall below these predicted quantiles, $\hat{\tau}_{j}=\frac{1}{N_{\text{test}}}\sum_{i=1}^{N_{\text{test}}}\mathbb{I}\left(y_{i}\leq\hat{y}_{ij}\right)$, with the nominal level $\tau_{j}$, for $j=1,2,\cdots,m-1$. If the model perfectly estimates the $\tau_{j}$-quantile, $\hat{\tau}_{j}$ should be close to $\tau_{j}$, for $j=1,2,\cdots,m-1$. We conduct the experiment over 100 replications and plot the average $\hat{\tau}_{j}$ across these replications in Figure 6. The result demonstrates a close alignment between $\hat{\tau}_{j}$ and $\tau_{j}$ for all $j=1,2,\cdots,m-1$, confirming that the model provides accurate quantile estimates.
@@
Next, we evaluate the generative capabilities of the QRGMM model through both unconditional and conditional tests, comparing the distributions of generated data against the true test set. Our comparisons include not only the CWGAN baseline, but also the original linear version of QRGMM, in order to highlight the gains brought by integrating the DeepFM component into QRGMM. Additionally, to better understand the individual contributions of the deep neural network component and the factorization machine component in QRGMM, we also perform ablation studies with two simplified variants: QRGMM-D (with only the deep neural network component) and QRGMM-F (with only the factorization machine component).
For CWGAN, we follow the hyperparameter choices recommended in athey2021using, and conduct a grid search over the sensitive parameter max_epochs. For QRGMM, we choose $m=\sqrt{n}$, with the remaining hyperparameters handled as in ordinary neural network training. QRGMM, QRGMM-D, and QRGMM-F adopt the same choice of $m$, and all hyperparameters of QRGMM-D and QRGMM-F are kept identical to those of QRGMM; the sole distinction lies in the output layer in Figure 5, where only a single component output is passed into the output layer instead of concatenating both components.
@@
Using the trained models, we generate random observations $\hat{y}^{\prime}_{i}$ for each test covariate vector $\mathbf{x}_{i}^{\prime}$, thus constructing generated test datasets. Then we can compare the overall distribution of the generated test observations $\{\hat{y}^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$ with the true test observations $\{y^{\prime}_{i}:i=1,\ldots,N_{\text{test}}\}$. We conduct 100 experimental replications, and calculate the sample means and standard deviations for the generated datasets and the true test dataset in each run. The averages of these values across the 100 replications are reported in Table 1. To quantify distributional similarity, we also computed the Wasserstein distance (WD) and the Kolmogorov-Smirnov statistic (KS) between the generated datasets and the true test dataset for each replication. The histograms of these measures across replications are illustrated in Figures 7 and 8.
@@
The above comparison is unconditional, assessing overall distributional similarity. Since we know the true data-generating process in the synthetic setting, we can also perform a conditional test to evaluate the generative performance conditioned on a specific covariate vector. For a given covariate vector $\mathbf{x}$, we generate conditional datasets of 10,000 observations of $\hat{Y}(\mathbf{x})$ using trained models and compare them with the true conditional dataset of 10,000 observations of $Y(\mathbf{x})$ sampled from the true model. Like the unconditional test, we perform 100 replications, calculating the sample means and standard deviations for each conditional dataset. The averaged results are also included in Table 1. The WD and KS values for the conditional test across replications are shown in Figures 9 and 10.
Table 1: Mean and standard deviation (SD) of generated observations.
@@
Truth $4355.4783$ $1430.8676$ $3948.9500$ $150.6615$
QRGMM $4350.6580$ $1412.5260$ $3948.7998$ $150.1433$
QRGMM $4355.9730$ $1421.0893$ $4114.5597$ $221.1148$
QRGMM-D $216.7717$ $117.6981$ $205.5038$ $96.5551$
QRGMM-F $4351.5033$ $1412.2491$ $3950.5061$ $150.8685$
CWGAN $4320.6589$ $1394.2039$ $3915.1583$ $190.8393$
Figure 7: WD for unconditional test.
Figure 8: KS for unconditional test.
Figure 9: WD for conditional test.
Figure 10: KS for conditional test.
@@
Figure 11: Histograms of generated vs. true data for unconditional (upper) and conditional (lower) test.
@@
Additionally, we visualize the performance of generative models by presenting histograms of the generated and true data distributions from one replication in Figure 11 for both the unconditional test and conditional test.
Summary of the Test Results
@@
Our results demonstrate that QRGMM consistently achieves significantly lower WD and KS values compared to CWGAN in both unconditional and conditional tests (see Table 1 and Figures 710). This indicates that QRGMM provides a more accurate and robust reproduction of the true data-generating process, whereas CWGAN exhibits considerable variability in performance across different replications. Furthermore, the histograms in Figure 11 illustrate that QRGMM-generated observations closely match the true data distribution in both unconditional and conditional settings, whereas CWGAN-generated observations exhibit noticeable discrepancies in conditional settings.
Although the original linear QRGMM attains slightly better mean and standard deviation alignment in the unconditional test, WD/KS values and visual plots reveal that its learned distribution is less accurate than QRGMM. In conditional tests, QRGMM clearly outperforms all other methods, highlighting its reliability in decision-making contexts where conditioning is essential. The results also demonstrate that it is essential to incorporate the DeepFM component in the model.
The ablation studies further shed light on the role of the deep neural network and factorization machine components. The results suggest that the factorization machine component plays a crucial role in enabling the model to quickly capture salient structural information. While the deep neural network component is more expressive and in principle subsumes the factorization machine component, using it alone makes it difficult for the model to learn accurate distributions. QRGMM-F performs much better than QRGMM-D but still falls short of the full QRGMM. Overall, the results highlight the benefit of incorporating DeepFM into QRGMM, while underscoring its robustness and fidelity in capturing complex data distributions.
6.1.2 Risk Measures Estimation
@@
In this synthetic-data experiment, for a specified $\mathbf{x}$, the data-generating process of $Y(\mathbf{x})$ is known. It enables us to analytically compute true risk measures as benchmarks for assessing the performance of QRGMM estimators. In our setting, the loss function $L(l)$ defined in Equation (1) is used, and the maximum loan amount $\bar{l}$ is set to the 99.9th percentile of $rY(\mathbf{x})$, with net unit revenue $r=1$. Loan levels $l$ are evaluated at 100 equally spaced points across $2000,\bar{l}]$.
@@
We compute four true risk measures: $r_{1}(l)$ (probability of default), $r_{2}(l)$ (expected loss), and $\text{LGD}(l)$ (loss given default), as specified in Section 2.2, and $r_{3}(l)=\mathbb{E}\left[\left(\left(l-rY(\mathbf{x})\right)^{+}\right)^{2}\right]$, a specific form of the generalized risk measure highlighting severe losses. QRGMM, QRGMM, and CWGAN estimators are calculated over 100 replications, each with 10,000 observations, and their means are compared to the true theoretical values, as well as the means of the estimators using the true test data observations calculated over 100 replications. These results are depicted in Figure 12, with subplots for four risk measures $r_{1}(l)$, $r_{2}(l)$, $\text{LGD}(l)$, and $r_{3}(l)$.
Figure 12: Estimation of risk measures across loan levels $l$.
The results are highly encouraging: the QRGMM estimators consistently track the true values across all risk measures, demonstrating robust accuracy. By contrast, QRGMM and CWGAN estimators show noticeable discrepancies overall, indicating weaker performance. This comparison underscores QRGMM’s superior capability in estimating complex risk measures, demonstrating its applicability in SCF credit risk management.
Furthermore, Figure 12 also offers a clear, immediate view of how risk measures evolve with loan amount $l$, providing decision-makers with an intuitive tool for understanding the risk dynamics, thereby greatly facilitating loan sizing decisions.
The Online Appendix C examines additional conditional tests at different covariate points. The results confirm that changes in covariates lead to noticeable shifts in conditional distributions and loan sizing outcomes, while QRGMM estimators remain stable and accurate across these variations.
@@ 6.2.1 Big Mart Sales Dataset
Figure 13: Alignment of nominal and empirical quantile levels on the Big Mart sales data.
Unconditional Distribution Comparison
Since the true data-generating mechanism is unknown in real-world datasets, we cannot perform a conditional test as in the synthetic experiment. Instead, we focus on an unconditional comparison of the generated and observed sales distributions. We trained the models on the training set and subsequently constructed a generated test dataset for each model across 100 replications.
Table 2 summarizes the means and standard deviations (averaged over the 100 replications) of both the generated and observed test sales. QRGMM and QRGMM show closer matches to the true data than CWGAN. To further quantify distributional similarity, we compute the WD and KS values for each replication, and plot their distributions in Figures 14 and 15. QRGMM again exhibits lower WD and KS values than CWGAN and QRGMM, confirming its superior generative ability in this real-world scenario.
Consistent with the experiments in Section 6.1, the results here reaffirm that the factorization machine component is essential while the deep neural network component provides complementary refinement, and their integration in QRGMM delivers the most accurate generative performance.
Table 2: Mean and standard deviation of generated observations.
@@
   Truth    $2177.0654$    $1708.2110$
   QRGMM    $2168.6467$    $1682.5005$
   QRGMM    $2173.0349$    $1637.7796$
   QRGMM-D    $6.7772$    $3.9158$
   QRGMM-F    $2168.0144$    $1681.5131$
   CWGAN    $2202.2440$    $1647.7403$
Figure 14: WD for Big Mart sales.
Figure 15: KS for Big Mart sales.
Figure 16 illustrates the observed and generated sales distributions in one representative replication. The distribution generated by QRGMM aligns more closely with the empirical distribution than those generated by CWGAN and QRGMM.
Figure 16: Histogram comparison of observed vs. generated Big Mart sales (unconditional test).
Risk Measures Estimation
For the Big Mart sales dataset, the conditional data-generating process of $Y(\mathbf{x})$ is not directly available, which prevents us from analytically computing the true conditional risk measures. To address this, we compute risk measures using the unconditional distribution of the original sales data and use them as benchmarks. Specifically, we evaluate the same four risk measures, $r_{1}(l)$, $r_{2}(l)$, $\text{LGD}(l)$, and $r_{3}(l)$, as described in Section 6.1.2, using $\bar{l}$ as the 99th percentile of sales, and compare the estimators from QRGMM, QRGMM, and CWGAN with the benchmarks.
The results, reported in Figure 17, show that QRGMM overall provides the closest approximations to the empirical benchmarks across all measures, consistently outperforming QRGMM and CWGAN. In particular, the estimators of $r_{1}(l)$ and $\text{LGD}(l)$ from QRGMM and CWGAN deteriorate when $l<2000$, as these models fail to adequately capture the left tail of the sales distribution (see also Figure 16). By contrast, QRGMM maintains strong performance in this region and achieves superior fidelity across the entire range, highlighting its robustness and accuracy in estimating risk measures for real-world sales data.
Figure 17: Estimation of risk measures across loan levels $l$.
Discussion
In summary, QRGMM demonstrates consistently strong performance on the Big Mart sales dataset, excelling in quantile estimation, distribution generation, and risk measures estimation. Although a conditional test is infeasible due to the unknown data-generating mechanism, unconditional evaluations, including distributional measures, visual diagnostics, and risk measure benchmarks, consistently show QRGMM outperforming both QRGMM and CWGAN. These findings highlight the practicality of QRGMM for complex SCF applications.
@@ 6.2.2 Amazon CBEC Dataset
Unconditional Distribution Comparison
Since the true conditional distribution remains unknown, we again adopt an unconditional evaluation strategy. All models are trained on the training set and used to construct generated test sales data over 100 replications.
Table 3 reports the means and standard deviations (averaged across 100 replications) of the generated and observed test sales, while Figures 19 and 20 compare the WD and KS values across replications. Overall, QRGMM continues to perform strongly, aligning closely with the true data and consistently yielding lower WD and KS values than CWGAN. In this case, QRGMM also performs reasonably well, indicating that the basic version remains effective in data-limited settings. These results confirm the robustness of QRGMM across different data conditions, while providing empirical evidence that the simpler QRGMM can also be competitive in practice when the sample size is limited.
Consistent with the synthetic and Big Mart results, the results on the Amazon CBEC dataset further confirm that the factorization machine is critical, which provides efficient second-order feature interaction learning in SCF applications, while the deep neural network provides complementary refinement, and their integration in QRGMM achieves the best performance.
Table 3: Mean and standard deviation of generated observations.
@@
  QRGMM   $39.3147$   $76.7319$
  QRGMM   $43.3484$   $94.6634$
  QRGMM-D   $13.8928$   $29.8903$
  QRGMM-F   $39.2903$   $76.7562$
  CWGAN   $39.8529$   $76.0303$
Figure 19: WD for Amazon CBEC sales.
Figure 20: KS for Amazon CBEC sales.
Figure 21 presents a histogram comparison of observed and generated sales distributions from one representative replication. As with the Big Mart dataset, the QRGMM-generated observations align more closely with the observed sales, whereas CWGAN shows noticeable deviation.
Figure 21: Histogram comparison of observed vs. generated Amazon CBEC sales (unconditional test).
Risk Measures Estimation
Similar to the Big Mart sales dataset, we also compute risk measures for the Amazon CBEC dataset using the unconditional distribution of the observed sales data and use them as benchmarks. As illustrated in Figure 22, overall, QRGMM appears to yield closer approximations to the empirical benchmarks, with performance generally exceeding those of QRGMM and CWGAN. The performance gap is most evident when $l<100$, where QRGMM and CWGAN fail to capture the left tail of the sales distribution, whereas QRGMM preserves stable and accurate estimation.
Figure 22: Estimation of risk measures across loan levels $l$.
Discussion
Overall, QRGMM continues to demonstrate robust performance on the Amazon CBEC dataset across quantile estimation, unconditional distribution generation, and risk measures estimation. Despite the relatively small sample size and the complex interaction effects among product selection teams and sales teams, QRGMM effectively captures the underlying data structure and provides stable estimation across all measures, underscoring its practical utility in complex CBEC environments and validating its applicability to SCF risk management.
Remark 5.
From a deployment perspective, the proposed framework can be integrated into existing credit assessment pipelines with only minor adjustments, as it primarily requires generative models of sales distributions that can be coupled with the standard risk evaluation modules already employed by financial institutions. Moreover, the data requirements are realistic within the SCF context, and the datasets used in our experiments reflect practical availability. On the one hand, lenders can draw upon large-scale historical sales datasets from e-commerce platforms such as Amazon, whether accumulated through prior lending experience, dedicated collection efforts, or licensed purchases from data providers. On the other hand, individual small- and medium-sized sellers applying for financing are typically required to authorize access to their sales and settlement information through official platform APIs (e.g., the Amazon Selling Partner API), thereby ensuring the availability of borrower-specific information and real-time performance data. This dual channel of data availability supports both robust calibration of generative models and tailored credit assessments. Potential regulatory implications primarily concern data privacy, cross-border information sharing, and model transparency. However, these challenges can be systematically managed through secure API protocols, repayment-account control mechanisms, and adherence to established financial data governance standards. Taken together, these considerations suggest that the proposed framework is not only technically deployable but also aligned with the broader evolution of data-driven, inclusive finance.